OpenStudy (anonymous):

A company distributes free erasers to all the students of x schools. Each school has (x + 2) classes. The number of students in each class is 1 less than the number of classes in each school. Each student is given (x - 2) erasers. Part A: Write an expression to show the total number of erasers distributed by the company in x schools. Part B: What would x(x + 2) represent? When simplified, what would be the degree and classification of this expression? Part C: How can you calculate the total number of students in each school?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

number of school =x number of classes of each school=(x+2) total number of classes=x(x+2)

OpenStudy (anonymous):

number of student= (x-1) total number of students= x(x+2)(x-1)

OpenStudy (anonymous):

i hope that you got the idea....

OpenStudy (anonymous):

So would I multiply x(x+2)(x-1) by (x-2)?

OpenStudy (anonymous):

i have something to do now but i will be back yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got x^4 -x^3 - 4x^2 +4x

OpenStudy (anonymous):

it looks good